OpenStudy (anonymous):

f(x)=g(x) f(x)=2x^2-x+1 g(x)=x^2-6x+8 Solve the Function.

OpenStudy (anonymous):

There are 2 answers.

OpenStudy (turingtest):

\[2x^2-x+1=x^2-6x+8\]\[x^2+5x-7=0\]now use the quadratic formula, which gives TWO answers for x because of the "plus or minus" part of the formula. You remember it, right?

OpenStudy (turingtest):

\[x=(-b \pm \sqrt{b^2-4ac})/2a\]in this case we have a=1 b=5 c=-7 plug in and solve!

OpenStudy (anonymous):

I understand that but how did you get to that point. Can you show all of your work please?

OpenStudy (turingtest):

I thought I did except I didn't first write f(x)=g(x) from there on it's what I have above: 2x^2-x+1=x^2-6x+8 x^2+5x-7=0 then quadratic and we're done.

OpenStudy (anonymous):

oh i was plugging in one of those for x. i see what i was doing wrong. THANKS!

OpenStudy (turingtest):

anytime!

OpenStudy (anonymous):

:)